a. What is the pH of 16.13 L solution if 19.89 g of sodium hydroxide are dissolved inside?

b. What is the pH of a 0.045 M HCl solution?
Answer #1

(Answer a) Given volume of Naoh solution 16.13 L weight of Naoh dissolved 19.899m. Gram. molecular weight of Nadt = 40gm conc7log (OH-] not [OH-] 3.08x10-2 Mele litt РОН a - log (3-08X10-2] 3.08 = 2-log 2-0.488 lo poh loslu pH + polt = 14 auton PH 14PH - logo lo(H+) 4.5 X10-2 -log 10 4.5 2-log 2-0.6532 PH = 1.346 The pH value of 03045 M HU 1:34 Scanned with CamScanner